Consider the M/M/16 queuing system λ=8 μ=14 and p = λ/(sμ)

L = L_{q}+\frac{\lambda }{\mu }

L_{q}=P(J\geq s)(\frac{\rho }{1-\rho })

PU > s) = (s!)(1-p)

W_{q}=\frac{L_{q}}{\lambda }

(a) average number of customers in the system

(b) average waiting time of each customer who enters the system

(c) probability that all servers are occupied



PU > s) = (s!)(1-p)
Answer #1

Using the formulas

PU > s) = (s!)(1-p)

L_{q}=P(J\geq s)(\frac{\rho }{1-\rho })

L = L_{q}+\frac{\lambda }{\mu }

a)Ls = 0.5714
b) Ws = 0.0714
c)
P(J >= 16) = 0